Evaluate the following project with the help of ERR for external
reinvestment rate 20% per year and MARR 18% per year. 7
Initial investment =Rs. 4,000,000
Expected life =8 Years
Salvage value =Rs. 50,000
Annual expenses =Rs. 160,000
Annual revenue =Rs. 320,000
1
Expert's answer
2020-05-26T11:21:21-0400

"NPV1=-4,000,000+\frac{320,000-160,000}{1+0.2}+\frac{160,000}{(1+0.2)^2}+\frac{160,000}{(1+0.2)^3}+\frac{160,000}{(1+0.2)^4}+\frac{160,000}{(1+0.2)^5}+\frac{160,000}{(1+0.2)^6}+\frac{160,000}{(1+0.2)^7}+\frac{160,000-50,000}{(1+0.2)^8}=-3,397,682.83"

"NPV2=4,000,000*(\frac{1.18}{1.2})^8=3,496,762.06"

"NPV1=3,397,682.83"

"NPV2=3,496,762.06"

"NPV1<NPV2" project is not effective